# Economic DAO Governance: A Contestable Control Approach

### v3

@@ 6 Concluding Thoughts
We have shown that there is a plausible mechanism for DAO governance that is code feasible, EV-robust, and efficient in the sense that it favors the business plan with the highest social surplus. The auction aspect of the mechanism provides a way out of the potential indeterminacy and pathologies that arise from pure voting mechanisms. At the same time, the sequential aspects of the mechanism create the ability to secure and enhance value that stems from DAO voting and other non-market governance approaches. The fact that empty voting is irrelevant to prevailing under the mechanism also removes the possible arbitrariness or deliberate harm that might result. The form of the mechanism permits the possible resolution of some major regulatory dilemmas. The mechanism tends to minimize the added token value received by project creators, thereby tending to maximize the investment value of the project, albeit imperfectly. In the rest of this section, we consider the mechanism in light of web3 and DAO governance norms.
@@
It is important to note that the auction mechanism does not ban conventional voting or vote buying regimes such as quadratic voting.114114114As discussed in Lalley and Weyl (2018), when quadratic voting operates well, it aggregates preferences and information accurately with respect to decisions. It is doubtful that quadratic voting could play a strong role in the context here in the face of asymmetric information and the possibility of free-riding. Parties without adequate information about potential control parties will not be able to represent their own interests well by voting, and revelation of information will trigger free-riding. If a project creator expends a large amount in a quadratic vote, the project creator will not be able to cover it from added token value due to that free-riding. If the DAO returns amounts the creator paid for the votes required to prevail to the creator, voting incentives are distorted.
@@
These approaches and the ensuing wisdom of crowds or attractive aggregative properties would operate fully during the open periods interspersed among control periods. Furthermore, even during control periods, the control party can abstain partially or fully on particular votes in order to allow the aggregative benefits of voting to be realized based on the control party’s calculus of how much expertise it has on those votes compared to the mass of other voters.115115115See Bar-Isaac and Shapiro (2020). Some DAO projects already conform to a rigid version of this more flexible possible control period behavior. A foundation or other centralized element implicitly or explicitly makes certain decisions with respect to the DAO while others are left entirely up the decentralized DAO governance mechanism. The auction mechanism here could extend to some current projects as a unified whole, encompassing both the centralized and decentralized portions of the projects. This approach would create on-going flexibility to allocate decisions between voting mechanisms and more centralized methods in the overall enterprise by using control periods to restructure the balance between the on-going centralized and decentralized elements. This ability to renew and redesign the structure in a coherent way would be perpetually available.
@@
That perpetual ability to renew and redesign the unified enterprise also addresses another rigidity, one that is intertemporal. Many enterprises begin with centralized incubation of the DAO element and then reach a point where governance passes entirely to a “fully decentralized” state, relying only on the DAO governance mechanism.116116116MakerDAO is a prominent example. Maker Foundation was a centralized entity used to develop the MakerDAO system. As described in MakerDAO (2020), the Maker Protocol Whitepaper: The Maker Foundation currently plays a role, along with independent actors, in maintaining the Maker Protocol and expanding its usage worldwide, while facilitating Governance. However, the Maker Foundation plans to dissolve once MakerDAO can manage Governance completely on its own. In July 2021, MakerDAO reached a point described by the Maker Foundation CEO as being “completely decentralized,” with the transfer of all Foundation functions to the DAO and the subsequent envisioned formal liquidation of the Foundation. Christensen (2021). One potential serious problem with this approach is that the enterprise may benefit at a later point from a major reorganization or revamp, perhaps with innovation in mind, that is critical to maintaining on-going dynamism or even for survival of the project.117117117Exactly this situation appears to have emerged for Maker. Only a short time after it presumably became completely decentralized, see note 116 supra, Rune Christensen, one of the co-founders, used his large token position to initiate a major reorganization of Maker through a project called “Endgame.” The goal was to bring Maker to the point where it could operate at a mass market level. Because of Christensen’s dominant token position, implementation of Endgame came about thorough an effective re-centralization of the project, a step that Christensen envisioned as temporary. See Gilbert, Aleks (2023). This re-organization or revamp may be difficult or impossible to accomplish without reinstating a later period of centralization that incubates a rebirth of the DAO in a renewed and more valuable form.118118118Relevant in this regard is Vitalik Buterin’s lament concerning Ethereum in 2022. Buterin (2022). Buterin stated that he had “diminishing influence” over Ethereum and that it was “becoming harder to make big changes to the Ethereum protocol due to the many stakeholders that have a say in the decision-making process,” creating an Ethereum that was “definitely more vetocratic.” He stated that: “Even now, I feel like the window is closing on substantial things. It’s getting harder to do big things even today.” Ethereum is not governed by a DAO, but the possibility of getting to a stale point is equally applicable to DAOs. Reorganization and major revamps are frequent in the corporate and non-profit world. It is not clear why DAOs or Ethereum would be any different with respect to the possible benefits of occasional major overhauls. It is very hard to see why restricting centralization to the beginning of the project is optimal in general. The auction mechanism leaves open the possibility of major reform during a control period by enabling consistent and coherent execution of a new plan similar to the plans used to birth DAOs during an initial period of centralized incubation. A control period for major reform can be invoked through the mechanism on a when-needed basis, with the most promising plan the odds-on favorite to be implemented.
@@ A.1 Proofs
Suppose that the bidder in fact can make the highest break-even bid $A_{1}^{*}$ of any of the auction participants and that the second highest break-even bid is $A_{2}^{*}$. Then the bidder can capture $\psi_{e}=\min\{(1-t_{m})\psi^{*},A_{1}^{*}-A_{2}^{*}\}$ of additional social surplus from the business plan $(V^{*},C^{*})$. The bidder accomplishes that by adding $\psi_{e}$ to $R$ in the bid as if it were an additional cost, which will reduce the bid by $\psi_{e}$.119119119The bidder can increase the resulting reduced bid by a small amount $\epsilon$ if there is a need to defeat a competing bid of exactly $A_{2}^{*}$. Retaining the business plan $(V^{*},C^{*})$ and adjusting $A_{1}^{*}$ as indicated dominates changing the business plan and submitting a different bid based on the alternative plan. Because the alternative plan will result in total social surplus $\psi_{a}<\psi^{*}$, a bid based on the alternative plan that directs the same amount of social surplus to the bidder will be lower than the corresponding bid under the original plan.
@@ A.2 Hedging and Stochastic Adjustments
The same expected value relationship does not apply for the value deposit and surety deposit. These deposits function as derivatives positions under the associated DAO Code rules. The value deposit is analogous to a bearish put spread, the analogy being exact in the case of the standard token deposit forfeit function with no baseline loss penalty.120120120See supra note 32 and the accompanying text. With respect to outcomes below $P_{0}$, a large put position with appropriate strike prices would be required to offset potential forfeits involving both the value deposit and surety deposit. The $T_{1}$ token holders, who benefit if the deposits are forfeited in part or entirely by the control party, effectively hold all these derivatives positions. When the control party succeeds in bringing the token value up to $V^{*}$ in the deterministic setting, both deposits are fully refunded and the $T_{1}$ token holders receive nothing. In a stochastic setting, $V^{*}$ is an expected value, not a certainty. There is a probability that the token price will never reach $V^{*}$ due to stochastic elements unrelated to the cogency of control party’s business plan and the effectiveness of execution of the plan. In that case, the control party will lose part or all of the value deposit, and if the stochastic elements are strong enough on the downside to result in a token price below $P_{0}$, part or all of the surety deposit. There is no corresponding upside: The deposits are fully refunded if $V^{*}$ is reached on a sustained basis, but attaining a higher level than $V^{*}$ results in no additional benefit. The overall result is that in the stochastic case, the implicit derivatives positions inherent in the two deposits both have a negative expected value for the winning bidder, while in the deterministic case they had zero expected value.

### v4

@@ 6 Concluding Thoughts
We have shown that there is a plausible mechanism for DAO governance that is code feasible, EV-robust, and efficient in the sense that it favors the business plan with the highest social surplus. The auction aspect of the mechanism provides a way out of the potential indeterminacy and pathologies that arise from pure voting mechanisms. At the same time, the sequential aspects of the mechanism create the ability to secure and enhance value that stems from DAO voting and other non-market governance approaches. The fact that empty voting is irrelevant to prevailing under the mechanism also removes the possible arbitrariness or deliberate harm that might result. The form of the mechanism permits the possible resolution of some major regulatory dilemmas and creates a strong case for a reduced level of regulation that is nonetheless at least equally effective. The mechanism tends to minimize the added token value received by project creators, thereby tending to maximize the investment value of the project, albeit imperfectly. In the rest of this section, we consider the mechanism in light of web3 and DAO governance norms.
@@
It is important to note that the auction mechanism does not ban conventional voting or vote buying regimes such as quadratic voting.120120120As discussed in Lalley and Weyl (2018), when quadratic voting operates well, it aggregates preferences and information accurately with respect to decisions. It is doubtful that quadratic voting could play a strong role in the context here in the face of asymmetric information and the possibility of free-riding. Parties without adequate information about potential control parties will not be able to represent their own interests well by voting, and revelation of information will trigger free-riding. If a project creator expends a large amount in a quadratic vote, the project creator will not be able to cover it from added token value due to that free-riding. If the DAO returns amounts the creator paid for the votes required to prevail to the creator, voting incentives are distorted.
@@
These approaches and the ensuing wisdom of crowds or attractive aggregative properties would operate fully during the open periods interspersed among control periods. Furthermore, even during control periods, the control party can abstain partially or fully on particular votes in order to allow the aggregative benefits of voting to be realized based on the control party’s calculus of how much expertise it has on those votes compared to the mass of other voters.121121121See Bar-Isaac and Shapiro (2020). Some DAO projects already conform to a rigid version of this more flexible possible control period behavior. A foundation or other centralized element implicitly or explicitly makes certain decisions with respect to the DAO while others are left entirely up the decentralized DAO governance mechanism. The auction mechanism here could extend to some current projects as a unified whole, encompassing both the centralized and decentralized portions of the projects. This approach would create on-going flexibility to allocate decisions between voting mechanisms and more centralized methods in the overall enterprise by using control periods to restructure the balance between the on-going centralized and decentralized elements. This ability to renew and redesign the structure in a coherent way would be perpetually available.
@@
That perpetual ability to renew and redesign the unified enterprise also addresses another rigidity, one that is intertemporal. Many enterprises begin with centralized incubation of the DAO element and then reach a point where governance passes entirely to a “fully decentralized” state, relying only on the DAO governance mechanism.122122122MakerDAO is a prominent example. Maker Foundation was a centralized entity used to develop the MakerDAO system. As described in MakerDAO (2020), the Maker Protocol Whitepaper: The Maker Foundation currently plays a role, along with independent actors, in maintaining the Maker Protocol and expanding its usage worldwide, while facilitating Governance. However, the Maker Foundation plans to dissolve once MakerDAO can manage Governance completely on its own. In July 2021, MakerDAO reached a point described by the Maker Foundation CEO as being “completely decentralized,” with the transfer of all Foundation functions to the DAO and the subsequent envisioned formal liquidation of the Foundation. Christensen (2021). One potential serious problem with this approach is that the enterprise may benefit at a later point from a major reorganization or revamp, perhaps with innovation in mind, that is critical to maintaining on-going dynamism or even for survival of the project.123123123Exactly this situation appears to have emerged for MakerDAO. Only a short time after it presumably became completely decentralized, see note 122 supra, Rune Christensen, one of the co-founders, used his large token position to initiate a major reorganization of MakerDAO through a project called “Endgame.” The goal was to bring MakerDAO to the point where it could operate at a mass market level. Because of Christensen’s dominant token position, implementation of Endgame came about thorough an effective re-centralization of the project, a step that Christensen envisioned as temporary. See Gilbert, Aleks (2023). This re-organization or revamp may be difficult or impossible to accomplish without reinstating a later period of centralization that incubates a rebirth of the DAO in a renewed and more valuable form.124124124Relevant in this regard is Vitalik Buterin’s lament concerning Ethereum in 2022. Buterin (2022). Buterin stated that he had “diminishing influence” over Ethereum and that it was “becoming harder to make big changes to the Ethereum protocol due to the many stakeholders that have a say in the decision-making process,” creating an Ethereum that was “definitely more vetocratic.” He stated that: “Even now, I feel like the window is closing on substantial things. It’s getting harder to do big things even today.” Ethereum is not governed by a DAO, but the possibility of getting to a stale point is equally applicable to DAOs. Reorganization and major revamps are frequent in the corporate and non-profit world. It is not clear why DAOs or Ethereum would be any different with respect to the possible benefits of occasional major overhauls. It is very hard to see why restricting centralization to the beginning of the project is optimal in general. The auction mechanism leaves open the possibility of major reform during a control period by enabling consistent and coherent execution of a new plan similar to the plans used to birth DAOs during an initial period of centralized incubation. A control period for major reform can be invoked through the mechanism on a when-needed basis, with the most promising plan the odds-on favorite to be implemented.
@@ A.1 Proofs
Suppose that the bidder in fact can make the highest break-even bid $A_{1}^{*}$ of any of the auction participants and that the second highest break-even bid is $A_{2}^{*}$. Then the bidder can capture $\psi_{e}=\min\{(1-t_{m})\psi^{*},A_{1}^{*}-A_{2}^{*}\}$ of additional social surplus from the business plan $(V^{*},C^{*})$. The bidder accomplishes that by adding $\psi_{e}$ to $R$ in the bid as if it were an additional cost, which will reduce the bid by $\psi_{e}$.125125125The bidder can increase the resulting reduced bid by a small amount $\epsilon$ if there is a need to defeat a competing bid of exactly $A_{2}^{*}$. Retaining the business plan $(V^{*},C^{*})$ and adjusting $A_{1}^{*}$ as indicated dominates changing the business plan and submitting a different bid based on the alternative plan. Because the alternative plan will result in total social surplus $\psi_{a}<\psi^{*}$, a bid based on the alternative plan that directs the same amount of social surplus to the bidder will be lower than the corresponding bid under the original plan.
@@ A.2 Hedging and Stochastic Adjustments
The same expected value relationship does not apply for the value deposit and surety deposit. These deposits function as derivatives positions under the associated DAO Code rules. The value deposit is analogous to a bearish put spread, the analogy being exact in the case of the standard token deposit forfeit function with no baseline loss penalty.126126126See supra note 31 and the accompanying text. With respect to outcomes below $P_{0}$, a large put position with appropriate strike prices would be required to offset potential forfeits involving both the value deposit and surety deposit. The $T_{1}$ token holders, who benefit if the deposits are forfeited in part or entirely by the control party, effectively hold all these derivatives positions. When the control party succeeds in bringing the token value up to $V^{*}$ in the deterministic setting, both deposits are fully refunded and the $T_{1}$ token holders receive nothing. In a stochastic setting, $V^{*}$ is an expected value, not a certainty. There is a probability that the token price will never reach $V^{*}$ due to stochastic elements unrelated to the cogency of control party’s business plan and the effectiveness of execution of the plan. In that case, the control party will lose part or all of the value deposit, and if the stochastic elements are strong enough on the downside to result in a token price below $P_{0}$, part or all of the surety deposit. There is no corresponding upside: The deposits are fully refunded if $V^{*}$ is reached on a sustained basis, but attaining a higher level than $V^{*}$ results in no additional benefit. The overall result is that in the stochastic case, the implicit derivatives positions inherent in the two deposits both have a negative expected value for the winning bidder, while in the deterministic case they had zero expected value.